# Attention as an Adaptive Filter

### v2

@@ Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By imposing a linear dynamics model with simultaneously diagonalizable state matrices and noise covariances, we can make use of a closed-form solution of the differential Lyapunov equation to efficiently propagate pairwise uncertainties through the dynamics. Attention naturally arises as the maximum likelihood solution for this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated pairwise precisions. To obtain a simplified variant with the same computational and memory complexity as standard attention, we further constrain the state matrix’s eigenvalues by tying their real parts to a shared scalar, i.e. enforcing isotropic decay. In the limit of vanishing dynamics and process noise, and using a small-angle approximation, we recover a complex-valued form of ordinary dot-product attention with rotary positional encodings.
@@ 1 Introduction
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence. Unlike SSMs, which are typically formulated as deterministic ODEs with inputs entering explicitly through control terms, we instead treat the entire dataset as arising from a stochastic differential equation (SDE), i.e. an autonomous system driven by noise.
@@
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and the closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously diagonalizable process and measurement noise covariances. Finally, we discuss an adaptively reweighted MLE that reweights propagated covariances using the pair-wise residuals.
@@
We generalize the robustified MLE to a tensor form of attention, and briefly analyze its time and memory complexity. We show that the algorithm can be simplified to a form with the same time and memory complexity as standard attention, through the use of convolutional kernels for the (diagonalized) matrix exponential and pair-wise propagated covariances, and by imposing additional structure on the linear SDE. We show that, in the case of a skew-symmetric state matrix and normalized inputs, this reduces to a complex-valued variant of standard attention with a complex-valued positional encoding equivalent to that in xPos and RetNet. We then describe an implementation of this algorithm, and an unrolled, recursive approximation for inference. We describe basic experiments on simple simulated linear dynamical systems.
@@
We generalize the SDE model from Section 3, which assumed isotropic noise, to allow for separate noise covariances in the tangential and radial directions (which we call a radial-tangential SDE). We show that the propagated covariances in this case can be written in closed-form since the assumed diagonal plus rank-1 covariances co-rotate with the dynamics. This formulation naturally yields a maximum likelihood estimator (MLE) on the hypersphere, and a corresponding attention mechanism. From this perspective, the Transformer can be interpreted geometrically, as an approximate solution of the MLE under the constraints imposed by the radial-tangential SDE model.
@@
We overview related work, including classical methods for filtering and adaptive filtering; the use of neural networks for modeling and filtering dynamical systems; neural state space models; connections between attention and structured state space models; related architectures; and various other topics.
@@
Appendix A: Background Additional details for the material in the preliminaries section.
@@
Extensions of AFA, a recurrent formulation, generative modeling, an extension to control, an extension to higher dimensional data, discussion of normal plus rank-1 noise covariance, and discussion of the expressivity of our structured SDE model.
@@ 2 Preliminaries
This section introduces the setup of our filtering problem and key ideas from linear Gaussian filtering on which we base the attention mechanism developed in Section 3. Detailed derivations can be found in Appendix A.
@@ 2.2 Batch Maximum Likelihood Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $\mathbf{x}(t_{i})$.
To simplify the notation, let $\mathbf{x}_{i}=\mathbf{x}(t_{i})$, $\mathbf{\hat{x}}_{ij}=\mathbf{\hat{x}}_{i}(t_{j})$ and $\mathbf{P}_{ij}=\mathbf{P}(\Delta t_{ij})$, etc. The (causal) MLE solution for $\mathbf{z}^{C}(t_{i})$ is a precision-weighted estimator:
@@
2.3 Robust Maximum Likelihood Trajectory Estimation
@@
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual has distribution $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}\big)$. We generally can’t observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}+\mathbf{R})$. We can then rescale the precision matrices with an adaptive weight:
$\mathbf{P}^{C}_{ij}\leftarrow\mathbf{P}^{C}_{ij}/w_{ij},\quad\text{where: }\,w_{ij}=\frac{1}{1+\mathbf{r_{z}}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r_{z}}_{ij}/\nu},$
where $\nu\in\mathbb{R}$ is a scaling factor. This leads naturally to an iteratively-reweighted least squares (IRLS) iteration in the eigenbasis. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}=e^{-\bm{\Lambda}\Delta t_{ij}}$, and define a corresponding measurement residual in the eigenbasis $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$. Let $\bm{\lambda}_{\mathbf{P}ij}^{C}=\text{diag}(\bm{\Lambda}_{\mathbf{P}ij}^{C})$. We apply the iterative update:
$\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k+1)}=\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},\quad\text{where}\quad\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)}=\left(\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\right)^{-1}\odot\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{C}\odot\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},$
@@
$w_{ij}^{(k)}=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{C})^{\top}\big|\mathbf{r}^{(k)}_{\mathbf{s}ij}\big|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
@@
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. (Alternatively, if we only have a single iteration, we can replace $\bm{\lambda}_{\mathbf{P}ij}^{C}$ with $\bm{\lambda}_{\mathbf{P}ij}^{C}+\bm{\lambda}_{R}$ (the covariance of $\mathbf{z}_{\mathbf{s}i}$).) To stabilize the iteration, we can instead use an affine combination of the current and previous estimate (i.e. a residual connection):
$\mathbf{\hat{z}}^{(k+1)}_{\mathbf{s}ii}=(1-\delta)\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k)}+\delta\mathbf{\hat{z}}_{\mathbf{s}i}^{\text{MLE}(k)},$
@@
This IRLS iteration, with precision matrices derived from an LTI model, can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\hat{z}}^{(k)}_{ii}\rightarrow\mathbf{z}^{C}$ in expectation with increasing $k$, we can take $\mathbf{\bar{z}}_{\mathbf{s}}\approx\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}$ at the last layer, and then predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{+}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}$

### v3

@@ Abstract
We introduce Adaptive Filter Attention (AFA), a novel attention mechanism that incorporates a learnable dynamics model directly into the computation of attention weights. Rather than comparing queries and keys directly, we model the input sequence as discrete observations of a linear stochastic differential equation (SDE). By assuming a continuous-time linear time-invariant system with simultaneously-diagonalizable state matrices and noise covariances, we can make use of a closed-form solution of the differential Lyapunov equation to efficiently propagate uncertainties through the dynamics from keys to queries. A generalization of attention naturally arises as the maximum likelihood solution for filtering the trajectory of this linear SDE, with attention weights corresponding to robust residual-based reweightings of the propagated query-key precisions. We further constrain the system dynamics and noise in order to obtain a simplified variant with the same computational and memory complexity as standard attention. In the limit of zero decay and process noise, and using a small-angle approximation, we recover a complex-valued generalization of ordinary dot-product attention with rotary positional encodings.
@@ 1 Introduction
Our goal is to bridge these two paradigms. We view self-attention as a form of filtering: it selectively averages across inputs, denoising by pooling information from similar states. However, unlike Kalman Filtering, standard attention does not explicitly propagate states through time. This motivates incorporating explicit temporal structure—through known or learned dynamics—into attention, restoring the inductive bias that standard attention loses by forgoing recurrence. Unlike SSMs, which are typically formulated as deterministic ODEs with inputs entering explicitly through control terms, we model sequence data as arising from a stochastic differential equation (SDE), i.e., an autonomous system driven by noise.
More precisely, rather than assuming the SDE describes the true global data-generating process, we adopt it as a local model from the perspective of each token, as a tractable means of propagating states and covariances across the sequence. Under the assumption of a linear system with simultaneously-diagonalizable state matrix and noise covariances, the differential Lyapunov equation admits a closed-form solution. This allows us to efficiently compute the covariance matrices and their inverses (precision matrices), which in turn enables the batch maximum likelihood estimation (MLE) of the system’s trajectory to be solved analytically. Incorporating residual-based robust reweightings of the propagated precisions, and embedding this adaptive filter as a neural network layer, we obtain a generalization of the attention mechanism. Thus, attention can be rigorously interpreted as the adaptive filtering solution for a linear SDE.
@@
We introduce the model setup and review standard results for propagating covariances through linear SDEs. We then discuss the batch maximum likelihood estimator and the closed-form solution of the differential Lyapunov equation for diagonalizable systems with simultaneously-diagonalizable process and measurement noise covariances. Finally, we discuss a robust estimator that reweights propagated covariances using the pair-wise residuals.
@@
We generalize the robustified MLE to a tensor form of attention, Adaptive Filter Attention (AFA), and analyze its time and memory complexity. We show that AFA can be simplified to a form with the same time and memory complexity as standard attention by assuming an isotropic-plus-skew-symmetric state matrix, and process and measurement noise covariances that are scalar multiples of identity, and we discuss the advantages of this architecture, which we call Isotropic AFA. We show that, in the case of a skew-symmetric state matrix and normalized inputs, this reduces to a complex-valued variant of standard attention with a complex-valued positional encoding that generalizes rotary positional encodings.
Section 4: Implementation
We describe an implementation of Isotropic AFA, including efficient computation of complex-valued operations. We then describe basic experiments on simple simulated linear dynamical systems.
@@
We generalize the LTI-SDE model from Section 3, which assumed isotropic process and measurement noise, to allow for separate noise covariances in the tangential and radial directions, which we call a "radial-tangential" SDE. We show that the propagated covariances in this case can be written in closed-form since the assumed diagonal plus rank-1 covariances co-rotate with the dynamics. This formulation naturally yields a maximum likelihood estimator (MLE) on the hypersphere, and a corresponding attention mechanism. From this perspective, the Transformer can be interpreted geometrically, as an approximate solution of the MLE under the constraints imposed by the radial-tangential SDE model.
@@
We overview related work, including classical methods for filtering and adaptive filtering; the use of neural networks for modeling and filtering dynamical systems; neural state space models; connections between attention and structured state space models; rotary positional encodings; related architectures; and geometrical interpretations of Transformers.
@@
Appendix A: Background
Detailed derivations for the material in the preliminaries section.
@@
Extensions of AFA, including for generative modeling, a version using a separable approximation of the full precision kernel, a recurrent formulation, an extension to control, an extension to higher dimensional data, additional discussion of normal plus rank-1 noise covariance, and discussion of the expressivity of radial-tangential linear SDEs.
@@ 2 Preliminaries
This section introduces the setup of our filtering problem and key ideas from linear Gaussian filtering, on which we base the attention mechanism developed in Section 3. Detailed derivations can be found in Appendix A.
@@ 2.2 Batch Maximum Likelihood Estimation
We wish to estimate the trajectory of the system from $m$ measurements taken at times $\{t_{1},...,t_{m}\}$, which we denote $\{\mathbf{z}(t_{1}),...,\mathbf{z}(t_{m})\}$. We can do so by solving a separate maximum likelihood estimation (MLE) problem for each $t_{i}$.
To simplify the notation, let $\mathbf{z}_{i}=\mathbf{z}(t_{i})$, $\mathbf{\hat{z}}_{ij}=\mathbf{\hat{z}}_{i}(t_{j})$, $\mathbf{P}_{ij}=\mathbf{P}(\Delta t_{ij})$, etc. The (causal) MLE solution for $\mathbf{z}^{C}(t_{i})$ is a precision-weighted estimator:
@@
2.3 Robust Estimation
Maximum Likelihood Estimation (MLE) of a system’s trajectory relies on accurate precision matrices, which represent the uncertainty of our estimates. If our dynamic model, defined by $\{\mathbf{S},\mathbf{S}^{-1},\bm{\Lambda},\bm{\Lambda}_{C},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}\}$, does not perfectly match reality, these theoretical precision matrices may misrepresent the true uncertainty, leading to suboptimal state estimates.
@@ 2.3 Robust Estimation
One approach to mitigate model mis-specification is to adaptively adjust the precision matrix $\mathbf{P}^{C}(\Delta t_{ij})$ using observed residuals. Since $\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\left(\mathbf{z}^{C}_{i},\mathbf{V}^{C}_{ij}\right)$, it follows that the pair-wise residual has distribution $\mathbf{r}_{ij}:=\mathbf{z}_{i}^{C}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{C}_{ij}\big)$. We generally cannot observe $\mathbf{z}_{i}^{C}$, but we can also define a measurement residual, whose covariance is inflated by the measurement noise $\mathbf{R}_{\text{curr}}$ at point $i$: $\mathbf{r_{z}}_{ij}:=\mathbf{z}_{i}-\mathbf{\hat{z}}_{ij}\sim\mathcal{N}\big(0,\mathbf{V}^{Z}_{ij})$, where $\mathbf{V}^{Z}_{ij}:=\mathbf{V}^{C}_{ij}+\mathbf{R}_{\text{curr}}$. Hence, we will define an adjusted total covariance $\mathbf{V}^{Z}:=\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{Z}\mathbf{S}^{-1}$, where:
@@
$\bm{\Lambda}_{\mathbf{V}}^{Z}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R,\text{proc}}+\bm{\Lambda}_{R,\text{curr}}$
Remark: Although the LTI SDE model assumes time-invariant process and measurement noise, the AFA framework ultimately aims to function as an adaptive, locally optimal filter. Each MLE estimate $\mathbf{\bar{z}}_{i}$ is treated as an independent optimization centered at its anchor time $t_{i}$, rather than under a single global LTI constraint, which would require $\bm{\Lambda}_{R,\text{proc}}=\bm{\Lambda}_{R,\text{curr}}$. By treating the MLE at $t_{i}$ as a local optimization problem, we allow the total propagated covariance to incorporate two separate, learned components of measurement noise: $\bm{\Lambda}_{R,\text{proc}}$ models the noise of the past observations that are propagated forward, and $\bm{\Lambda}_{R,\text{curr}}$ models the static, intrinsic noise of the current anchor point $\mathbf{z}_{i}$. The linear dynamics are used only as a computationally efficient prior to propagate observations via the time-dependent kernel $\mathbf{V}_{C}(\Delta t_{ij})$, not as a global constraint on the MLEs for every point in the trajectory. The separate $\bm{\Lambda}_{R,\text{curr}}$ term acts as a global noise floor, which also gives the model the ability to decouple the propagated precision from the dynamics propagation, and to learn that the noise process is not perfectly time-invariant across the sequence.
We replace $\mathbf{P}^{C}_{ij}$ in the MLE with a robust reweighting based on the residuals.
$\mathbf{P}^{C}_{ij}\leftarrow w_{ij}\mathbf{P}^{Z}_{ij}$
where we have chosen a robust weight of the form:
$w_{ij}=\frac{1}{1+\mathbf{r_{z}}_{ij}^{\top}\mathbf{P}^{Z}_{ij}\mathbf{r_{z}}_{ij}/\nu},$
where $\nu\in\mathbb{R}$ is a scaling factor.
Here, $\mathbf{P}^{Z}_{ij}$ encodes the directional confidence from the LTI model, while $w_{ij}$ adaptively rescales the precision based on the observed residual $\mathbf{r}_{ij}$, down-weighting measurements with larger than expected deviations.
Remark: The term $\mathbf{r}^{\top}\mathbf{P}\mathbf{r}$ is known as the Mahalanobis distance, which quantifies the difference between a vector $\mathbf{r}$ and a distribution with precision $\mathbf{P}$. The similarity score $1/(1+\mathbf{r}^{\top}\mathbf{P}\mathbf{r}/\nu)$ behaves similarly to the unnormalized softmax: if one residual is much smaller than the others, it’s corresponding weight $w_{ij}$ dominates, while if all residuals are similar, the output approximates a uniform weighting. The key difference arises when all residuals are small relative to $\nu$: unlike softmax, which picks a dominant term whenever relative differences are large, this similarity score allows for a roughly uniform attention distribution when all of the residuals are small. In this sense, $\nu$ acts as a kind of statistical evidence threshold. Using the fact that $\log(1+x)\approx x,\,\,\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{Z}_{ij}\mathbf{r}_{ij}/\nu}\approx\exp(-\mathbf{r}_{ij}^{\top}\mathbf{P}^{Z}_{ij}\mathbf{r}_{ij}/\nu).$
2.4 Iteratively-Reweighted Least Squares
This robust-reweighting must take place in the eigenbasis. Let $\mathbf{z}_{\mathbf{s}i}:=\mathbf{S}^{-1}\mathbf{z}_{i}$ and $\mathbf{\hat{z}}_{\mathbf{s}ij}=\mathbf{S}^{-1}\mathbf{\hat{z}}_{ij}$, and define a corresponding measurement residual in the eigenbasis $\mathbf{r_{zs}}_{ij}:=\mathbf{z}_{\mathbf{s}i}-\mathbf{\hat{z}}_{\mathbf{s}ij}$.
Letting $\bm{\lambda}_{\mathbf{P}ij}^{Z}=\text{diag}(\bm{\Lambda}_{\mathbf{P}ij}^{Z})$, and applying a robust-reweighting of the diagonalized MLE results in a single step in an iteratively-reweighted least squares (IRLS) iteration in the eigenbasis:
$\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k+1)}=\mathbf{\bar{z}}_{\mathbf{s}i}^{(k)},\quad\text{where}\quad\mathbf{\bar{z}}_{\mathbf{s}i}^{(k)}=\left(\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{Z}\right)^{-1}\odot\sum_{j}w_{ij}^{(k)}\bm{\lambda}_{\mathbf{P}ij}^{Z}\odot\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)},$
$w_{\mathbf{s}ij}^{(k)}:=\frac{1}{1+(\bm{\lambda}_{\mathbf{P}ij}^{Z})^{\top}\big|\mathbf{r}^{(k)}_{\mathbf{s}ij}\big|^{2}/\nu},\quad\text{where}\quad\mathbf{r}_{\mathbf{s}ij}^{(k)}:=\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ii}-\mathbf{\hat{z}}^{(k)}_{\mathbf{s}ij}$.
@@ 2.4 Iteratively-Reweighted Least Squares
and initial condition $\mathbf{\hat{z}}_{\mathbf{s}ii}^{(1)}=\mathbf{z}_{\mathbf{s}i}$. To stabilize the iteration, we can use an affine combination of the current and previous estimate (i.e. a residual connection):
$\mathbf{\hat{z}}^{(k+1)}_{\mathbf{s}ii}=(1-\delta)\mathbf{\hat{z}}_{\mathbf{s}ii}^{(k)}+\delta\mathbf{\bar{z}}_{\mathbf{s}i}^{(k)},$
@@
This IRLS iteration, with precision matrices derived from an LTI model, can be implemented with a multi-layer neural network with learnable parameters $\mathbf{S},\bm{\Lambda},\bm{\Lambda}_{Q},\bm{\Lambda}_{R}$ shared across layers. If $\mathbf{\bar{z}}^{(k)}_{\mathbf{s}i}\rightarrow\mathbf{z}^{C}_{\mathbf{s}i}$ in expectation with increasing $k$, we can take our estimate at the last layer $\mathbf{\bar{z}}^{(k)}_{\mathbf{s}i}$ and use it to predict the measurement at the next time step in the eigenbasis using the (diagonalized) matrix exponential:
$\mathbf{z}^{+}_{\mathbf{s}i}=\bm{\Lambda}_{C}e^{\bm{\Lambda}(t_{i+1}-t_{i})}\bm{\Lambda}_{C}^{-1}\mathbf{\bar{z}}_{\mathbf{s}i}^{(k)}=e^{\bm{\Lambda}(t_{i+1}-t_{i})}\mathbf{\bar{z}}_{\mathbf{s}i}^{(k)}$
@@
Hence, the same learned dynamics model governs uncertainty propagation (covariance $\mathbf{V}(\Delta t_{ij})$), state estimation (propagating past measurements to the current anchor $\mathbf{\hat{z}}_{\mathbf{s}ij}$), and future prediction (advancing the final estimate $\mathbf{\bar{z}}_{si}^{(k)}$ forward in time). Enforcing this self-consistency across uncertainty modeling, filtering, and prediction compels the model to learn a dynamics model (parameterized by $\bm{\Lambda}$) that jointly optimizes all three processes, which is a form of regularization to encourage temporal consistency.